OpenStudy (anonymous):

A right circular cylinder has a base whose area is 616 square feet. If the height of the cylinder is 13 feet, what is its lateral surface area? Use π = Answer 363 ft2 572 ft2 1144 ft2 1760 ft2

OpenStudy (anonymous):

Area of base = 616 = πr2 => r=sqrt(616/π)=14 Circle length = 2πr = 2π14 = 88 Lateral are = height*circle length = 88*13 = 1144